Fermi National Laboratory


2005 Accelerator Update Archive

December 19, 2005-December 21, 2005
- During the 48 hour period operations established two store that along with a previous store provided the experiments with 38 hours and 30 minutes of luminosity.

December 16, 2005-December 19, 2005
- During the 72 hour period operations established three stores that provided the experiments with 56 hours and 30 minutes of luminosity.

December 14, 2005-December 16, 2005
- Separator spark causes TeV quench
- Pbar LCW leak found and fixed

December 12, 2005-December 14, 2005
- Two stores provided five hours and 14 minutes of luminosity
- TeV F0 ground fault found and corrected
- TeV suffers 3 quenches
- Booster LCW leak found and repaired
- TeV F17 collimator problem

December 9, 2005-December 12, 2005
- Tevatron separators conditioning
- NuMI off due to horn chiller problems
- Pbar began stacking
- Recycler began stashing
- TeV put in shot setup, but suffered quench

December 7, 2005-December 9, 2005
- TeV has VCB problem
- Pbar Debuncher has power supply problem
- Switchyard has power supply problem
- First store may go into the TeV late on Sunday

December 5, 2005-December 7, 2005
- NuMI hits integrated total of 1E20 protons on target
- CUB LCW leak
- TeV sector A4 cryo system helium leak repaired
- TeV sector B1 cryo system cold

December 2, 2005-December 5, 2005
- The TeV B1 mechanical and cryo system sign offs should occur by noon on Monday
- The TeV B1 cooldown will start shortly after the sign offs
- The TeV B1 separator bakeout should finish on Tuesday

November 30, 2005-December 2, 2005
- TeV B1 work continues
- TeV B11 separator bake out will extend TeV shutdown
- Pbar conducts reverse proton studies
- Recycler conducts Electron Cooling studies

November 28, 2005-November 30, 2005
- Tevatron sector B17 work continues
- MiniBooNE, NuMI, and the Meson experiments resume taking beam

November 23, 2005-November 28, 2005
- No stores established due to Tevatron B17 work
- Pbar's lithium lens fails
- Experts report that the B17 bypass is okay

November 21, 2005-November 23, 2005
- One store provided 8 hours and 6 minutes of luminosity
- Multi-house quench lost store 4515
- TeV sector B1 cryo system warming up

November 18, 2005-November 21, 2005
- Three stores provided 47 hours and 24 minutes of luminosity
- MiniBooNE has power supply problems
- P1 transfer line vacuum bursts
- Store 4515 quenched
- Pbar off due to LCW leak

November 16, 2005-November 18, 2005
- One store provided one hour and seven minutes of luminosity
- Accelerator safety system checks completed
- Linac RF stations repaired
- TeV vacuum and power supply problems
- Feeder 37 fault
- Store delayed due to Recycler RF structure problem

November 14, 2005-November 16, 2005
- Two stores provided 41 hours and 15 minutes of luminosity
- Safety system testing
- Ground water in MI-40
- Stack dumped to fix AP2 vacuum

November 11, 2005-November 14, 2005
- Three stores provided 62 hours and 23 minutes of luminosity
- Linac off for five hours due to tripped Klystron
- Recycler loses CE504 but holds on to stash
- MiniBooNE off due to H866

November 9, 2005-November 11, 2005
- Two stores provided 41 hours and 11 minutes of luminosity
- NuMI problems
- Electron Cooling back in action
- Store 4495 sets New Luminosity Record: 166.8E30

November 7, 2005-November 9, 2005
- Two stores provided 37 hours of luminosity to the experiments
- TeV RF spark caused quench
- Pelletron down for work
- Multiple problems troubles NuMI

November 4, 2005-November 7, 2005
- 4 stores provide 53 hours and 39 minutes of luminosity
- P1 suffers from vacuum bursts
- Recycler stash aborted

November 2, 2005-November 4, 2005
- Two stores provided 27 hours and 48 minutes of luminosity
- Store 4481 aborted due to RF trip

October 31, 2005-November 2, 2005
- Two stores provided 34 hours and 38 minutes of luminosity
- Store 4477 set a Record Luminosity of 164.17E30
- ICW leak at CDF
- Store 4477 quenched on termination

October 28, 2005-October 31, 2005
- During this 72 hour period: three stores provided 60 hours and 25 minutes of luminosity
- TeV quench, but with no store

October 26, 2005-October 28, 2005
- During this 48 hour period: One store provided 29 hours and 8 minutes of luminosity
- Store 4473 set a New Luminosity Record with a luminosity of 158.223E30

October 24, 2005-October 26, 2005
- During this 48 hour period: two stores provided 27 hours and 19 minutes of luminosity
- P1 vacuum bursts continue
- Store 4468 sets New Luminosity Record with an initial luminosity of 144.906E30
- TeV B2 VCB opened and caused store 4468 to abort
- D0 must open its detector for power supply work

October 21, 2005-October 24, 2005
- During this 72 hour period: Three stores provided 62 hours and 18 minutes of luminosity
- P1 vacuum cascades controlled
- Store 4465 sets Record Luminosity with an initial luminosity of 144.512E30

October 19, 2005-October 21, 2005
- During this 48 hour period: One store provided 34 hours and 38 minutes of luminosity
- P150 Vacuum Bursts
- TEL required timing adjustment
- D:IB trouble
- TeV crate and sequencer trouble

October 17, 2005-October 19, 2005
- During the 48 hour period: Two stores provided 20 hours and 22 minutes of luminosity
- Vacuum bursts plague TeV, Recycler, and MI
- Pbar pulse magnet ground faulted
- ComEd power glitch aborts store and stash
- TeV vacuum leak hunted

October 14, 2005-October 17, 2005
- During this 72 hour period: 4 stores provided 57 hours and 19 minutes of luminosity
- Meson LCW Leak
- TeV C1 Pump Fails
- Vacuum Failure Aborts Store
- Pelletron Off
- NuMI has Bad Chiller Compressor

October 12, 2005-October 14, 2005
- During this 48 hour period: Two stores provided 38 hours and 9 minutes of luminosity
- NuMI resumes taking beam
- Pelletron repaired
- Store 4445 sets Record Luminosity: 142.51E30

October 10, 2005-October 12, 2005
- During this 48 hour period: Two stores provided 15 hours and 40 minutes of luminosity
- CDF solenoid dumped helium
- Pelletron down
- Booster LCW repaired, then had ORBMP vacuum problem

October 7, 2005-October 10, 2005
- During this 72 hour period: one store provided 29 hours and 38 minutes of luminosity
- Booster LCW leak

October 5, 2005-October 7, 2005
- NuMI horn groundfault cleared
- Close call for Tevatron Vacuum

October 3, 2005-October 5, 2005
- During this 48 hour period: two stores provided 37 hours and 52 minutes of luminosity
- Store 4431 sets New World Hadron Luminosity Record: 141.28E30

September 30, 2005-October 3, 2005
- During this 72 hour period: two stores provided 30 hours and 29 minutes of luminsoity
- NuMI Horn ground faulted
- Two TeV quenches during shot setup
- Recycler stash lost
- TeV separator problems

September 28, 2005-September 30, 2005
- During this 48 hour period: 1 store provided 33 hours and 45 minutes of luminosity
- TeV helium leak
- Pbar AP1 vacuum leak
- NuMI Horn problems

September 26, 2005-September 28, 2005
- During this 48 hour period: 1 store plus existing store provided 33 hours and 45 minutes of luminosity
- Recycler only store 4402: 119E30
- Pbar stack lost
- Two TeV quenches
- Recycler record for store antiprotons: 247E10

September 23, 2005-September 26, 2005
- During this 72 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 65 hours and 56 minutes of luminosity
- Pbar stacking problems
- Store 4401 sets New Luminosity Record - 131.02E30

September 21, 2005-September 23, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 44 hours and 34 minutes of luminosity
- NuMI kicker problems
- Stacking suffers from DRF1 troubles

September 19, 2005-September 21, 2005
- During this forty-eight hour period operations established one store that combined with an existing store provided the experiments with approximately 32 hours and 36 minutes of luminosity
- NuMI off due to dripping water
- Preaccelerator switched to the I- Source
- TeV quench

September 16, 2005-September 19, 2005
- During this 72 hour period operations established two stores that combined with an existing store provided approximately 58 hours and 39 minutes of luminosity to the experiments
- H- Source and 750 KeV line problems
- MI and Recycler RF trip due to LCW pump
- Pbar loses small stack

September 14, 2005-September 16, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 41 hours and 39 minutes of luminosity
- Pbar losses on Recycler stash
- H- Source problems

September 12, 2005-September 14, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 45 hours and 22 minutes of luminosity
- ComEd glitch causes loss of stash

September 9, 2005-September 12, 2005
- During this 72 hour period Operations established two stores that provided the experiments with approximately 51 hours and 27 minutes of luminosity
- Power glitch quenches store, and dumps stack and stash
- Pbar off due to D:QF failure

September 7, 2005-September 9, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 14 hours and 54 minutes of luminosity
- Meson suffers from LCW trouble
- Store luminosity ruined by separator spark
- Recycler loses antiprotons

September 5, 2005-September 7, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 40 hours of luminosity
- Booster EAPS problems
- M05 LCW leak

September 2, 2005-September 5, 2005
- During the 72 hour period operations established three stores that combined with an existing store provided the experiments with 41 hours of luminosity
- A bad vacuum card causes the loss of a Tevatron store

August 31, 2005-September 2, 2005
- Maintenance work was done on a number of machines
- PBar had some magnets moved as part of the maintenance period. This caused problems for stacking when PBar started back up
- After studies, tuning, and accesses, stacking returned to an acceptable level
- The machines were glitched and a store lost when Feeder 47 faulted again. Store 4363 was put in the Tevatron and was colliding this morning
- The experiments ran when beam was available

August 29, 2005-August 31, 2005
- During this forty eight hour period Operations established one store that combined with an existing store provided the experiments with approximately 44 hours and 15 minutes of luminosity
- New Pbar stach record for the Recycler Ring is 191.0 E10

August 26, 2005-August 29, 2005
- During the 72 hour period operations established three stores that combined with an existing store provided the experiments with 33 hours and 30 minutes of luminosity
- Two feeder trips and associated glitches make for a problematic weekend

August 24, 2005-August 26, 2005
- During the 48 hour period operations established one store that combined with an existing store provided the experiments with 42 hours and 22 minutes of luminosity
- Electron cooling down for Pelletron repairs
- NuMI HV101 troubles due to LCW leak

August 22, 2005-August 24, 2005
- During this forty eight hour period Operations established one store that combined with an existing store provided the experiments with approximately 44 hours and 58 minutes of luminosity
- MCenter held off due to LCW trouble

August 19, 2005-August 22, 2005
- During this 72 hour period Operations established three stores that combined with an existing store provided the experiments with approximately 39 hours of luminosity
- Pbar emittance problems
- Thunderstorm aborts store and quenches TeV
- Recycler stash lost

August 17, 2005-August 19, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with 35 hours and 7 minutes of luminosity
- Lost shots to Recycler investigated
- Lightning strike aborts store 4337

August 15, 2005-August 17, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with 22 hours and 40 minutes of luminosity
- Lithium lens repaired
- TeV quench
- F sector vacuum problems

August 12, 2005-August 15, 2005
- During this 72 hour period Operations established two stores that combined with an existing store provided approximately 69 hours and 6 minutes of luminosity to the experiments
- Pbar's lithium lens tripped off

August 10, 2005-August 12, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 27 hours and 35 minutes of luminosity
- MI RF causes NuMI troubles
- TeV antiproton kicker fails
- Recycler looses RF structure

August 8, 2005-August 10, 2005
- During this 48 hour period Operations established one store that when combined with an existing store provided the experiments with approximately 38 hours and 18 minutes of luminosity
- Booster RF limps through to Tuesday maintenance

August 5, 2005-August 8, 2005
- During this 72 hour period operations established two stores that combined with an existing store provided approximately 66 hours and 46 minutes of luminosity to the experiments
- Massive Booster RF problems

August 3, 2005-August 5, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 35 hours and 49 minutes of luminosity
- Store 4318 sets Luminosity Record with 129.8E30

August 1, 2005-August 3, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 41 hours and 26 minutes of luminosity
- Pelletron network problems

July 29, 2005-August 1, 2005
- During this 48 hour period operations established two stores that combined with an existing store that provided approximately 58 hours and 20 minutes of luminosity to the experiments
- Booster kicker keeps beam off
- TeV quench

July 27, 2005-July 29, 2005
- During this 48 hour period operation established two store that combined with an existing store provided the experiments with approximately 26 hours and 27 minutes of luminosity
- Stores lost due to quench and an inadvertent abort

July 25, 2005-July 27, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 28 hours and 10 minutes of luminosity
- MI & Booster repair RF systems
- TeV quench

July 22, 2005-July 25, 2005
- During the 72 hour period Operations established three store that provided the experiments with approximately 52 hours and 58 minutes of luminosity
- Feeder 47 trips off
- MI and Booster suffer from RF problems
- Antiproton Target stops rotating

July 20, 2005-July 22, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 25 hours and 19 minutes of luminosity
- MiniBooNE has DAQ trouble
- Feeder 47 tripped off
- TeV aborts store 4287
- TeV quench due to power supply trip

July 18, 2005-July 20, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 27 hours and 46 minutes of luminosity
- MI suffers from RF and damper problems
- RF trip causes TeV store abort

July 15, 2005-July 18, 2005
- During this 72 hour period Operations established two store that provided the experiments with approximately 64 hours and 45 minutes of luminosity

July 13, 2005-July 15, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 11 hours and 48 minutes of luminosity
- TeV experts conduct studies
- TeV quench looses store

July 11, 2005-July 13, 2005
- During this 48 hour period Operations established one store that provided the experiments with approximately 18 hours and 15 minutes of luminosity
- LRF4 conditioning slow
- Pbar troubles continue
- TeV lonitudinal damper problem

July 8, 2005-July 11, 2005
- During this 72 hour period Operations established one store that combined with an existing store provided the experiments with approximately 50 hours and 7 minutes of luminosity
- LRF4 conditioning slow
- Pbar troubled with many problems, ending with a ground fault
- TeV quench during shot setup
- Recycler showed first proof of electron cooling

July 6, 2005-July 8, 2005
- During this 48 hour period Operations established two store that combined with an existing store provided the experiments with 32 hours and 30 minutes of luminosity
- MI damper, beam loading, and feed forward problems
- New LRF4 PA has trouble conditioning

July 4, 2005-July 6, 2005
- During this 48 hour period Operations established one store that provided the experiments with approximately 25 hours and 51 minutes of luminosity
- NuMI off due to power supply
- MI transformer replaced Monday evening
- Store established Tuesday morning

July 1, 2005-July 4, 2005
- During this 72 hour period operations established one store that provide the experiments with approximately 29 hours and 46 minutes of luminosity
- Linac switched to I- Source
- KRF7 tripped off
- MI transformer ground fault

June 29, 2005-July 1, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided approximately 38 hours and 15 minutes of luminosity to the experiments

June 27, 2005-June 29, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 27 hours and 3 minutes of luminosity
- Heat trips off MiniBooNE beam
- TeV quench

June 24, 2005-June 27, 2005
- During this 72 hour period operations established three stores that provided the experiments with approximately 61 hours and 2 minutes of luminosity
- Booster RF sparking
- Low Beta Magnet movement lowers luminosity

June 22, 2005-June 24, 2005
- During this 48 hour period operations established one store that combined with an existing store provided the experiments with approximately 11 hours and 36 minutes of luminosity
- TeV suffers two quenches
- I- Source problems
- Accumulator ground fault
- TeV suffers abort - store lost
- Feeder fault
- 400 MeV chopper trip

June 20, 2005-June 22, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided approx. 40 hours and 13 minutes of luminosity to the experiments
- CUB LCW pump failed
- Mysterious "glitch" trips off TEL

June 17, 2005-June 20, 2005
- During this 72 hour period Operations established four stores that provided 46 hour and 6 minutes of luminosity to the experiments
- TeV quench due to wet engine-store lost
- B0 device causes store abort
- Recycler stash lost

June 15, 2005-June 17, 2005
- Operations established no stores during this 48 hour period
- Problems slowed shot setup
- Tev quenches due to separator spark

June 13, 2005-June 15, 2005
- During this 48 hour period Operations established no stores
- No NuMI magnet ground fault
- Pbar stack lost
- Feeder 38 and 47 reenergized
- Pbar Debuncher Cryo system ready

June 10, 2005-June 13, 2005
- During the past 72 hours Operations established 2 stores that provided approximately 46 hours and 27 minutes of luminosity to the experiments
- Booster chopper trips off
- I- Source voltage troubles
- CHL losses power; store terminated
- Possible bad NuMI magnet required MI access; stash terminated
- TeV beam not expected before evening shift June 15

June 8, 2005-June 10, 2005
- During this 48 hour period Operations established one store that provided the experiments with approximately 28 hours and 44 minutes of luminosity
- Booster and NuMI suffer from kicker problems
- Pbar stacking halted due to A:IB contactor
- TeV quench during shot setup

June 6, 2005-June 8, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 44 hours and 32 minutes of luminosity
- NuMI off four hours
- Booster off due to GMPS failure

June 3, 2005-June 6, 2005
- During the past 72 hour period Operations established two stores that provided the experiments with approximately 36 hours and 7 minutes of luminosity
- Store delayed by kicker problem and quench
- Pbar accesses to look for leak
- MCenter shutdown for five days of work

June 1, 2005-June 3, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 41 hours and 48 minutes of luminosity
- Booster chopper and notcher problems
- Power supply holds off NuMI

May 27, 2005-June 1, 2005
- During this 144 hour period Operations established 4 stores that combined with an existing store provide the experiments with approximately 101 hours and 41 minutes of luminosity
- Booster off for LCW leak and kicker spark
- Store aborted due to controller problem
- Debuncher Pulse Magnet replaced

May 25, 2005-May 27, 2005
- During this 48 hour period Operations established two stores that provided approximately 43 hours and 45 minutes of luminosity
- Many necessary repairs were done during this 48 hour period.

May 23, 2005-May 25, 2005
- During this 48 hour period Operations established two stores that provided the experiments with approximately 29 hours and 45 minutes of luminosity
- Problems with RF and Timing, along with a bad Wet engine flywheel make for long shot setups.

May 20, 2005-May 23, 2005
- During this 72 hour period Operations established two stores that provided the experiments with approximately 38 hours of luminosity
- TeV problems throughout the weekend

May 18, 2005-May 20, 2005
- During this 48 hour period Operations established three stores that provided approximately 8 hours and 3 minutes of luminosity
- TeV looses three stores due to three quenches

May 16, 2005-May 18, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 30 hours and 27 minutes of luminosity
- Booster suffered from RF, EAPS, and Chopper problems
- NuMI sent beam during Recycler shot setup
- TeV quenched after store 4148 terminated

May 13, 2005-May 16, 2005
- During this 72 hour period Operations established two stores that provided the experiments with approximately 52 hours and 1 minute of luminosity
- TeV quenches three times - one store lost
- I- Source problems
- The Booster Chopper trips

May 11, 2005-May 13, 2005
- During this 48 hour period operations established one store that combined with an existing store provided approximately 28 hours and 39 minutes of luminosity
- Store 4138 sets New Luminosity Record: 126.97E30
- Lightening causes loss of store, stack, and trips many systems
- MiniBooNE horn suffers cooling problem

May 9, 2005-May 11, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 42 hours and 9 minutes of luminosity
- Booster trips due to EAPS
- Pbar and NuMI suffer trips

May 6, 2005-May 9, 2005
- During this 72 hour period Operations established three stores that provided the experiments with approximately 62 hours and 27 minutes of luminosity
- Booster has chopper, RF, and power supply trouble
- Most of one recycler stash attempt lost in transfers
- NuMI target leak returns

May 4, 2005-May 6, 2005
- During this forty-eight hour period Operations established one store that provided the experiments with approximately 21 hours and 12 minutes of luminosity
- Safety system testing continues
- NuMI target problems return
- BRF4 removed for repairs

May 2, 2005-May 4, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided approximately 32 hours and 32 minutes of luminosity
- TeV quench loses store
- NuMI loses power

May 2, 2005-May 4, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided approximately 32 hours and 32 minutes of luminosity
- TeV quench loses store
- NuMI loses power

April 29, 2005-May 2, 2005
- During this 72 hour period Operation established three stores that provided the experiments with approximately 57 hours and 32 minutes of luminosity
- The TeV set a NEW RECORD Luminosity: 122.65E30
- TeV power supply, HP3US, springs LCW leak
- Pbar suffers from emittance problems
- Accumulator Water Skip fails
- NuMI has helium leak
- Two stores established with luminosities of 122E30

April 27, 2005-April 29, 2005
- During the 48 hour period Operations established one store that provided the experiments approximately 32 hours and 13 Minutes of luminosity
- Store 4116 set New Record with a luminosity of 119.03E30
- TeV suffers from C4 QPM and power supply problems

April 25, 2005-April 27, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 42 hours and 40 minutes of luminosity
- MiniBooNE resumes taking beam

April 22, 2005-April 25, 2005
- During this 72 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 62 hours and 35 minutes of luminosity
- C2 wet engine failure, store saved by quick action
- Booster kicker gets replaced
- Accumulator power supply trip looses stack

April 20, 2005-April 22, 2005
- During this 48 hour period operations established two stores that provided the experiments with approximately 31 hours and 43 minutes of luminosity
- Pbar septa interrupts stacking
- NuMI conducts target scan

April 18, 2005-April 20, 2005
- During this 48 hour period operations established one store that combined with an existing store provided the experiments with approximately 37 hours and 47 minutes of luminosity
- Main Injector suffers from RF cavity phase error
- Accumulator power supply (A:LQ) causes problems
- Accumulator Septa power supply causes problems

April 15, 2005-April 18, 2005
- During this 72 hour period operations established one store that provided approximately 25 hours and 41 minutes of luminosity to the the experiments
- A shorted P2 transfer line quadrupole halts stacking
- Magnet replacement completed Sunday morning

April 13, 2005-April 15, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 43 hours and 53 minutes of luminosity
- Stacking halted due to P1/P2 transfer line problem

April 11, 2005-April 13, 2005
- During the 48 hour period operations established two stores that combined with an existing store provided the experiments with approximately 27 hours and 58 minutes of luminosity
- TeV suffered two quenches
- Debuncher kicker fails

April 8, 2005-April 11, 2005
- During this 72 hour period Operations established two stores that provided the experiments with approximately 42 hours and 22 minutes of luminosity
- A Recycler only store quenched before it was established
- Booster had East Anode power supply problems

April 6, 2005-April 8, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments approximately 40 hours and 34 minutes of luminosity for the experiments
- Store 4078 quenched due to separator spark
- Recycler supplied antiprotons for store 4081
- Booster tunnel LCW leak

April 4, 2005-April 6, 2005
- During this 48 hour period Operations established two stores that provided the experiments with approximately 33 hours and 2 minutes of luminosity
- LRF5 PA tube replaced
- Booster chopper "on" tube replaced
- TeV quench during antiproton transfer

April 1, 2005-April 4, 2005
- During this 72 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 63 hours and 42 minutes of luminosity
- MiniBooNE magnet repaired
- Pbar and Recycler suffer from emittance problems
- TeV RF station trips off causing luminosity loss
- NuMI scans target

March 30, 2005-April 1, 2005
- During the last 48 hours Operations established two stores that combined with an existing store provided the experiments with approximately 12 hours and 17 minutes
- Off site power glitch causes store abort
- MiniBooNE magnet overheats

March 28, 2005-March 30, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 45 hours and 36 minutes of luminosity
- Booster suffers from RF problems
- Tevatron motion detectors record Sumatra earthquake
- Linac experts discover and repair RF station hose leaks
- I- Source motor generator failed
- Booster vacuum valves close
- Antiproton Source emittances affects stacking

March 25, 2005-March 28, 2005
- During this 72 hour period Operations established 3 stores that combined with an existing store provided approximately 62 hours and 20 minutes of luminosity to the experiments
- BRF16 proved troublesome all weekend
- KRF6 LCW leak caused minor problem

March 23, 2005-March 25, 2005
- During this 48 hour period Operations established two stores that, combined with an existing store, provided the experiments with approximately 41 hours and 52 minutes of luminosity
- Meson target train cooling trips off
- Pbar set new stacking record: 16.2mA per hour

March 21, 2005-March 23, 2005
- During this 48 hour period Operations established two stores that provided the experiments with approximately 38 hours and 15 minutes of luminosity
- Store 4053 set record luminosity: 117.32E30
- Store 4053 lost due to separator spark
- Booster has RF problems

March 18, 2005-March 21, 2005
- During this 72 hour period operations established two store that combined with an existing store provided the experiments with approximately 32 hours and 6 minutes of luminosity
- Store 4088 set record luminosity - 113.143E30
- Store 4088 quenched due to kicker prefire
- Kicker work and condition delays store
- B4 wet engine work delays study
- TeV suffers quench during study

March 16, 2005-March 18, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 30 hours and 36 minutes of luminosity
- D0 tested new luminosity counter
- NuMI intensity increasing

March 14, 2005-March 16, 2005
- During this 48 hour period operations established two stores that provided the experiments with approximately 41 hours and 33 minutes of luminosity
- Linac suffers RF and Buncher trips

March 11, 2005-March 14, 2005
- During this 72 hour period operations established two stores that provided the experiments with approximately 54 hours and 5 minutes of luminosity
- QPM system repaired
- NuMI losses beam due to LCW problem
- Operators save store 4041
- Antiproton Source has kicker problem
- Cryo technicians save store 4042

March 9, 2005-March 11, 2005
- During this 48 hour period Operations established one store that provided the experiments with approximately 2 hours and 37 minutes of luminosity
- Tevatron plagued with quenches and aborts

March 7, 2005-March 9, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 27 hours and 13 minutes of luminosity
- Pbar and NuMI commission new stacking TLG module
- KRF6 problems held off beam
- Booster Watchdog power supply held off beam
- H- Source suffers from intensity drop
- Store 4027 quenched
- Debuncher Kicker problem halted stacking
- TeV trim magnet fails

March 4, 2005-March 7, 2005
- During this 72 hour period, Operations established 3 stores that, combined with an existing store, provided the experiments with approximately 57 hours and 25 minutes of luminosity
- One TeV store lost due to quench
- NuMI experiment commissioned
- D-Zero solenoid quenches

March 2, 2005-March 4, 2005
- During this 48 hour period operations established one store, which provided the experiments with approximately 13 hours and 11 minutes of luminosity.
- Vacuum leak repaired in TeV
- TeV suffered quench in shot setup

February 28, 2005-March 2, 2005
- During the 48 hour period Operations established one store that provided approximately 29 hours and 15 minutes of luminosity for the experiments
- NuMI and Pbar take beam
- Booster has RF problems
- TeV quench losses store before it's established
- H- Source suffers arc current problem

February 25, 2005-February 28, 2005
- During this 72 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 52 hours and 39 minutes of luminosity.
- Booster has RF problems
- NuMI conducting studies
- Store 4007 quenched

February 23, 2005-February 25, 2005
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 40 hours and 8 minutes of luminosity
- Store 4002 sets NEW LUMINOSITY RECORD with a luminosity of 107.43E30
- MECAR problems hit the Main Injector
- MiniBooNE losses beam due to bad power supply

February 21, 2005-February 23, 2005
- During this 48 hour period, Operations established one store that combined with an existing store provided the experiments with approximately 37 hours and 35 minutes of luminosity
- Linac has quad power supply problems
- Main Injector has MECAR problems
- Recycler stashed antiprotons

February 18, 2005-February 21, 2005
- During this 72 hour period Operations established three stores that combined with an existing store provided approximately 29 hours and 23 minutes of luminosity to the experiments.
- Tevatron suffers three quenches
- Recycler losses stash
- NuMI begins taking beam

February 16, 2005-February 18, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 44 hours and 5 minutes of luminosity
- Recycler stashed antiprotons
- Booster has MP01 problem
- NuMI decay tunnel readied for beam

February 14, 2005-February 16, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 30 hours and 19 minutes of luminosity.
- Vacuum problems halt Pbar's ability to stack
- Booster suffers from RF problems
- Recycler losses stash
- Main Injector has vacuum problems

February 11, 2005-February 14, 2005
- During this 72 hour period, Operations established two stores that provided approximately 18 hours and 34 minutes of luminosity to the experiments
- Operations received permission to power the Main Injector and to run beam in all machines
- TeV separator and MI Septa problems delayed shots

February 9, 2005-February 11, 2005
- During this 48 hour period, Operations provided approximately 21 hours and 30 minutes of luminosity to the experiments.
- Main Injector work continues

February 7, 2005-February 9, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with approximately 26 hours and 19 minutes of luminosity.
- A quench ended store 3969
- The Main Injector held off due to ground fault indication
- MI-50 suffered a switch gear failure

February 4, 2005-February 7, 2005
- During this 72 hour period Operations established two stores that combined with an existing store delivered approximately 57 hours and 2 minutes of luminosity to the experiments
- A Tevatron quench lost store 3962
- Coyote helps troubleshoot MI Kicker problem
- A separator spark reduces store 3968's luminosity
- A Recycler sequencer problem limited store 3969 to Pbar antiprotons

February 2, 2005-February 4, 2005
- During this 48 hour period, operations established one store that, combined with an existing store, provided the experiments with approximately 25 hours and 30 minutes of luminosity.
- I:Q703 power supply replaced.

January 31, 2005-February 2, 2005
- During this 48 hour period Operations established two stores that provided approximately 32 hours and 49 minutes of luminosity to the experiments.
- A ComEd power glitch tripped off the TEL and ARF1, but store and stack survived.

January 28, 2005-January 31, 2005
- During the last 72 hours Operations established two stores that provided approximately 62 hours and 20 minutes of luminosity for the experiments.
- The Tevatron set a New Luminosity Record of 105.17E30 on Saturday evening.
- A Booster 4000 MeV chopper problem held off beam for a while.
- The motor generator for the I- Source had a brush problem.
- A problem with the G2 vacuum system brought down the Switchyard enclosure C system, which shut off beam to the Meson experiment.

January 26, 2005-January 28, 2005
- During this 48 hour period Operations established one store that combined with an existing store provided the experiments with 34 hours and 56 minutes of luminosity.
- MI-60 cubical removal completed on Wednesday morning.

January 24, 2005-January 26, 2005
- During this 48 hour period operations established one store that provided approximately 26 hours and 31 minutes of luminosity to the experiments.
- Booster experts have been working on the machine's tune
- The Antiproton Source has trouble with ARF1

January 21, 2005-January 24, 2005
- During this 72 hour period operations established one store that provided the experiments with approximately 4 hours and 36 minutes of luminosity
- Quenches, leaks, and wet engine problems plagued the Tevatron all weekend
- NuMI reported seeing their first neutrinos at the MINOS near detector on Friday

January 19, 2005-January 21, 2005
- During the last 48 hours operations established one store that, combined with an existing store, provided the experiments with approximately 19 hours and 25 minutes of luminosity
- The I- Source operations were troubled by a dirty water-resistor
- The TeV lost two stores due to aborts
- NTF treated its first patient on Thursday
- NuMI received permission to operate on Thursday. It will take its first beam on Friday

January 17, 2005-January 19, 2005
- During this 48 hour period, Operations established two stores that provided approximately 40 hours and 40 minutes of luminosity to the experiments.
- On Monday, the TeV set a New Luminosity Record of 104.96E30. Unfortunately, the next store was only 104.25E30.
- On Monday, the Antiproton Source set a New Stacking Record with a rate of 14.94 mA per hour.
- NTF has a problem with the power supply and magnet cabling for its 32-degree bend magnet.

January 14, 2005-January 17, 2005
- During this 72 hour period operations established 2 stores that, combined with an existing store, provided approximately 63 hours and 1 minute of luminosity to the experiments.
- A broken BPM has stopped the use of the Linac's 400 MeV steering program.
- Recycler stashed antiprotons and then used them for two stores
- A Tevatron separator trip caused a loss of luminosity

January 12, 2005-January 14, 2005
- During this 48 hour period, Operations established one store that, combined with an existing store, supplied the experiments with approximately 41 hours and 21 minutes of luminosity.
- A TeV BLM software problem held up shot setup
- A beam problem caused some CDF SVX crates to trip off
- A Booster kicker problem held off stacking
- A MI-60 VCB flash-over shutdown MI at 10 PM on Thursday night

January 10, 2005-January 12, 2005
- During this 48 hour period, Operations established one store that combined with an existing store provided approximately 44 hours and 41 minutes of luminosity to the experiments.
- The Recycler is regularly stashing antiprotons
- The Main Injector has a RF frequency problem
- The Antiproton Source achieves new stacking record

January 7, 2005-January 10, 2005
- During the last 72 hours, Operations established three stores that supplied the experiments with approximately 34 hours and 20 minutes of luminosity
- One store was lost due to an abort
- The Tevatron suffered two quenches that delayed shot setup
- Whenever possible, all stores will be made from Pbar and Recycler antiprotons

January 5, 2005-January 7, 2005
- Operations established one store during this 48 hour period. It provided approximately 26 hours and 39 minutes of luminosity to the experiments.
- LAM42 caused problems
- MBEX was also a source of problems

January 3, 2005-January 5, 2005
- Operations and the accelerator complex continued to recover from the ComEd power outage
- Tevatron used this down time to adjust the CDF low beta quad magnets
- Pbar resumed stacking on Tuesday afternoon, but then lost its stack that evening due to a breaker trip
- The TeV cryo system was ready by midnight on Tuesday

[an error occurred while processing this directive]

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.


Return to Current Accelerator Update



last modified 12/21/2005   email Fermilab